OpenStudy (anonymous):

What are the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1?

OpenStudy (anonymous):

you can follow this example Assuming the recursive definition is tn = 2*tn-1 t1 = 3 t2 = 2*t1 = 2*3 = 6 t3 = 2*t2 = 2*6 = 12 t4 = 2*t3 = 2*12 = 24

OpenStudy (anonymous):

ar^0+ar^1..... ar^n t1 t2 tn

OpenStudy (anonymous):

can you draw -3tn-1 for us please

OpenStudy (anonymous):

What do you mean?

OpenStudy (campbell_st):

is it \[t_{n} = -3t^{n -1}\]

OpenStudy (anonymous):

Yeah except the n-1 is not in exponential form...I don't know what it's called haha but I think you read it as -3t(sub)n-1 if that makes sense :p

OpenStudy (campbell_st):

ok... so is it \[t_{n} = -3t_{n-1}\]

OpenStudy (anonymous):

yupp

OpenStudy (campbell_st):

ok... that should help @timo86m

OpenStudy (campbell_st):

so the pattern is 4, 4*(-3), (4*-3)*-3, and (4*-3*-3)*-3

OpenStudy (campbell_st):

just evaluate for the terms

OpenStudy (campbell_st):

so the pattern is saying multiply -3 by the previous turn after starting at 4 ist term is 4 so the 2nd term is 4 times -3 3rd term is -3 times the 2nd term 4th term is -3 times the 3rd term

OpenStudy (anonymous):

Thank you! (: